OpenStudy (anonymous):

what is a way to turn 3/5 into a decimal besides dividing 3 by 5

OpenStudy (anonymous):

I'd multiply top and bottom by 2 \[\frac{ 3 }{ 5 } * \frac{ 2 }{ 2 } = \frac{ 6 }{ 10 } = 0.6\] we know what happens when divide by multiple of 10s divide by 10s= move numberator 1 left of the decimal pt.

OpenStudy (anonymous):

divide by 100s, move 2 left divide by 1000 move 3 lefts, etc.

OpenStudy (anonymous):

thank you for your help :)

OpenStudy (anonymous):

anytime :)